The revenue equation​ (in hunderds of millions of​ dollars) for barley production in a certain country is approximated by

R(x)=0.0632x2 + 1.2443x + 2.2473 where x is in hundreds of millions of bushels. Find the ​marginal-revenue equation and use it to find the marginal revenue for the production of the given number of bushels.

​(a) The​ marginal-revenue equation is R′(x)= -------------------------. ​(Round to four decimal places as​ needed.)

​(b) Find the marginal revenue for the production of 200,000,000 bushels.The marginal revenue is----------------

hundred million dollars. ​(Type an integer or decimal rounded to two decimal places as​ needed.)

​(c) Find the marginal revenue for the production of 550,000,000 bushels.The marginal revenue is----------------------- hundred million dollars. ​(Type an integer or decimal rounded to two decimal places as​ needed.)
Answer #1

2, + 2.243 2 d ( 063 2 12443x + 2, 24-73) :) _ R·(X)- 0.1264(x) + 112443 二 b) A、いく二 200,000,00 Mavml Miion(c) Since X= 550,000,000( in Hundred of million)

Marginal Revenue = (0.124)* 550,000,000 +1.2443

M.E(550,000,000)= 68,200,001.24 ( In Hundred of million)